Calculate Osp for calcium fluoride (Ksp = 3.9 x10-11) when 220.0 mL of a 6.70x10-3 M solution of Ca(NO3)2 is added to 225.0 mL of a 6.75x10-3 Msolution of KF.
Answer #1

Lets find the concentration after mixing for Ca(NO3)2

Concentration after mixing = mol of component / (total volume)

M(Ca(NO3)2) after mixing = M(Ca(NO3)2)*V(Ca(NO3)2)/(total volume)

M(Ca(NO3)2) after mixing = 0.0067 M*220.0 mL/(220.0+225.0)mL

M(Ca(NO3)2) after mixing = 3.312*10^-3 M

Lets find the concentration after mixing for KF

Concentration after mixing = mol of component / (total volume)

M(KF) after mixing = M(KF)*V(KF)/(total volume)

M(KF) after mixing = 0.00675 M*225.0 mL/(225.0+220.0)mL

M(KF) after mixing = 3.413*10^-3 M

So, we have now

[Ca2+] = 3.312*10^-3 M

[F-] = 3.413*10^-3 M

At equilibrium:

CaF2 <----> Ca2+ + 2 F-

Qsp = [Ca2+][F-]^2

Qsp = (3.312*10^-3)*(3.413*10^-3)^2

Qsp = 3.858*10^-8

we have,

Ksp = 3.9*10^-11

Since Qsp is greater than ksp, precipitate will form

Answer: Qsp = 3.9*10^-8